Find the length of a rectangle if the width is 3x-2 and tge perimeter is 20x+18
Answer by swincher4391(1107)

p = 2l + 2w
20x+18 = 2l + 2(3x-2)
20x+18 = 2l + 6x - 4
14x+22 = 2l
7x +11 = l